In a survey of a group of​ men, the heights in the​ 20-29 age group were normally​ distributed, with a mean of 67.3 inches and a standard deviation of 4.0 inches. A study participant is randomly selected. Complete parts​ (a) through​ (d) below. ​(a) Find the probability that a study participant has a height that is less than 68 inches. The probability that the study participant selected at random is less than 68 inches tall is nothing . ​(Round to four decimal places as​ needed.)
Answer #1

a)

Given,

\mu = 67.3 , \sigma = 4

We convert this to standard normal as

P( X < x) = P( Z < x - \mu / \sigma)

So,

P( X < 68) = P( Z < 68 - 67.3 / 4)

= P( Z < 0.175)

= 0.5695